40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 255 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  4523   Thu Apr 14 01:03:43 2011 KojiUpdateLSCLSC Campaign ~ Status

Target: To lock the Michelson with the new RF/LSC


Status

RF generation box: READY - already ready to go to the IOO rack. (Suresh)

RF distribution box: In Progress - the internal components are to be connected. (13th evening - Suresh)

Placing PD and CCD: Done - PD and CCD on the AP table (13th Afternoon - Aidan, Larisa with supervision of Kiwamu)

Cabling1: Done - PD signal AP table to the demodulator (13th Afternoon - Jamie with supervision of Suresh)

Cabling2: Done - RF generation box (IOO Rack) to the demodulator

Demodulator: In Progress - Test and install (13th night - Kiwamu with supervision of Suresh)

LSC model: Done - Run the new LSC model. (It is named as "C1LST" so far) (13th evening - Jamie)

LSC medm: Done on 14th - Modify the current LSC medm screens Update the EPICS database Adjust the matrices (- Jenne with supervision of Koji)

  4522   Thu Apr 14 00:21:28 2011 KojiUpdateCDSNew C1LSC code running

[Jamie, Jenne, Koji]

We installed the new c1lsc and started the process.

We still need to configure bunch of the EPICS variables, matrices, and some of the filters.
This should be done in order to transmit the signals to the suspensions.
Jenne is going to work on this task tomorrow (Friday) morning,
and Koji will take over the task afternoon/evening.

  4521   Wed Apr 13 23:32:07 2011 Aidan, JamieConfigurationLSCAS PD and Camera installed

I spent some time tracking down the AS beam which had vanished from the AP table. Eventually, by dramatically mis-aligning SRM, PRM and ITMY, returning BS to its Jan 1st PITCH and YAW values and tweaking the ITMX alignment [actual values to follow], I was able to get an AS beam out onto the AP table. I verified that it was the prompt reflection off ITMX by watching it move as I changed the YAW of that optic and watching it stay stationary as I changed the YAW of ITMY.

Jamie and I then steered the beam through a 2" PLCX-50.8-360.6 lens and placed the RF PD (AS55) at the focus. Additionally, we installed the AS camera to observe the leakage field through a Y1S steering mirror (as shown in the attached diagram).

Currently the PD has power but the RF and DC outputs are not connected to anything at the moment.

Atm 2 by Steve

 

 

Attachment 1: AS_beam.jpg
AS_beam.jpg
Attachment 2: P1070546.JPG
P1070546.JPG
  4520   Wed Apr 13 16:56:08 2011 BryanConfigurationGreen LockingY-ARM Green-Locked!

 Locked!

The Y-arm can now be locked with green light using the universal PDH servo. Modulation frequency is now 277kHz - chosen because it seems to produce smaller offsets due to AM effects

To lock, turn on the servo, align the system to give nice circular-looking TEM_00 resonances, and wait for a good one. It'll lock on a decent mode for a few seconds and then you can turn on the local boost and watch it lock for minutes and minutes and minutes.

The suspensions are bouncing around a bit on the Y-arm and the spot is quite low on the ETMY and a little low on ITMY, but from this point it can be tweaked and optimised.

 

 

 

  4519   Wed Apr 13 16:38:17 2011 Larisa ThorneUpdateElectronicsVideo MUX camera/monitor check

 [Kiwamu, Larisa]

 

The following Video MUX inputs(cameras) and outputs(monitors) have been checked:

MC2F, FI, AS Spare, ITMYF, ITMXF, ETMYF, ETMXF, PSL Spare, ETMXT, MC2T, POP, MC1F/MC3F, SRMF, ETMYT, PRM/BS, CRT1(MON1), ETMY Monitor, CRT2(MON2), CRT4(MON4), MC1 Monitor, CRT3(MON3), PSL1 Monitor, PSL2 Monitor, CRT6(MON6), CRT5(MON5), ETMX Monitor, MC2 Monitor, CRT9, CRT7(MON7), CRT10, and Projector.

 

Their respective statuses have been updated on the wiki:   (wiki is down at the moment, I will come back and add the link when it's back up)

  4518   Wed Apr 13 11:34:07 2011 josephbUpdateCDSFixed IFO_ALIGN.adl

Problem:

I switched the ITMX and ITMY control channels yesterday, but forgot to update the IFO_ALIGN.adl file (/opt/rtcds/caltech/c1/medm/c1ifo/) which had the control labels swapped to make life easier.

Solution:

I swapped ITMX and ITMY control locations on the screen.

Question:

Are there any other screens involving ITMX and ITMY that had controls reversed to make life easier?

  4517   Tue Apr 12 18:15:07 2011 kiwamuSummaryIOORF combiner is more like attenuator

I realized that my impedance matching theory on an RF combiner was wrong !

In fact an RF combiner behaves more like an attenuator according to a reflection measurement that I did today.

A 3-way combiner reduces power of an input signal by a factor of 4.8 dB because it can be also considered as a 3-way splitter.

So it is just a lossy component or in other words it is just an attenuator.

 

(reflection measurement)

To check my speculation that I posted on #4504 I measured reflection coefficients for both cases.

In the measurement I used a heliax cable, which goes from 1X2 rack to the PSL table with a length of about 10 m. Note that this is the cable that had been used as '33 MHz EOM'.

At the input of the heliax cable it was connected to a direction coupler to pick off reflections and the reflected signal was sampled in AG4395A.

The other end of the cable (output side of the cable) was basically connected to the resonant box.

Then I did a reflection measurement for both cases as drawn in this entry (see #4504).

  - case 1 -  the combiner was inserted at the input side of the heliax cable.

  - case 2 - the combiner was directly attached to the resonant box

On the combiner, ZFSC-3-13, the port 1 and 2 were terminated with 50 Ohm, therefore the port 3 was used as an input and the source port is the output.

Here is a resultant plot of the reflection measurements.

reflection_test.png

Note that whole data are calibrated so that it gives 0 dB when the output side of the heliax is open.

There are two things we can notice from this plot:

 (1) The reflection coefficient at the resonant frequencies (where notches appear) are the same for both cases.

 (2) Over the measured frequency range the reflections were attenuated by a factor of about 9.6 dB , which is twice as large as the insertion loss of the combiner.

These facts basically indicates that  the RF combiner behaves as a 4.8 dB attenuator.

Hence the location of the combiner doesn't change the situation in terms of RF reflections.

Quote from #4505

 The input impedance of the resonant box was measured when an RF combiner was attached to the box.

Indeed the combiner makes the impedance more 50 Ohm and reduces the reflection.

 

 

  4516   Tue Apr 12 16:01:33 2011 josephbUpdateGeneralRFM errors

Problem:

Currently the c1scy, c1mcs, and c1rfm models are reporting an error with receiving some data sent over the GE Fanuc Reflected memory cards.

To be more exact, the C1:SUS-ETMY_ALS signal from the c1gcv FE code on the c1ioo computer going too the Y end is not being received However, the C1:SUS-ETMY_LSC signal is.  So the physical RFM card seems to be working.

Similarly, the TRY signal is being sent correctly from the Y end computer.  The X end is working fine and receiving both LSC and ALS signals.

The c1mcs and c1rfm models also receive data from the c1ioo computer and reporting receiving errors.

Theory:

Because the RFM cards are transmitting and receiving at least some channels, I'm guessing there was changes made to the C1.ipc file, which defines the memory locations of these various channels on the RFM network, and that when a model was rebuilt, a different one using the previous IPC file was not, and thus one of the computer is going to the wrong place to either read or write data.

Tomorrow, I'm planning on the  following:

1) Clean out the C1.ipc file (/opt/rtcds/caltech/c1/chans/ipc/)

2) Rebuild all models

3) Run activate_daq.py script

4) Restart models via script

If this doesn't clear up the problem, I'll continue  to bug hunt.

  4515   Tue Apr 12 12:01:30 2011 josephbUpdateGeneralIFO controls, now with 10% less lying (ITMX/ITMY controls swapped)

The ITMX/ITMY control swap is complete.

The steps from this elog were followed.

In addition, I did a burt restore of c1sus, c1mcs.

I then swapped all the gain settings from ITMX to ITMY, and reenabled the watchdogs.

I did some basic kick tests (1000 counts into UL coil) and confirmed channels like C1:SUS-ITMX_ULPD_VAR (watchdogs mV readback) corresponded to the correct optic.  I also checked that the POS, PIT, YAW, SIDE produced reasonable damping when engaged.

  4514   Mon Apr 11 23:35:02 2011 ranaUpdateRF SystemInstalled low pass filters in the demod boards

I am a little concerned about using these low pass filters so close to the band edge. Recall that there is no on-board preamp for the RF input to the mixer.

So, if the input impedance of the filters is not 50 Ohms, we will get some unwanted reflections and sensitivity to cable length.

I think its worth while to check the impedance or S-parameters of these things with the LO activated to find out if we need to remove them or not.

  4513   Mon Apr 11 21:13:15 2011 KojiUpdateElectronicsNew Green PD test1

Correction:

The (-) input has been decoupled by the capacitor. So the series resistance of the PD is not the matter.
In this sense, we should use 0Ohm for the (+) input shunting.

Quote:

51 Ohm for CLC409

The datasheet of CLC409 uses 25Ohm there. This is to cancel the input bias current of the two inputs of the opamp.

The source impedance (series) of SGD444 is 50Ohm. So I used 50Ohm for the + input shunting.

However, I could probably use anything between 0-50Ohm as the datasheet itself tells that the bias currents are
not related between the two inputs. In addition, I am not sure how much the real series resistance of the PD is.

  4512   Mon Apr 11 20:03:05 2011 taraUpdateElectronicsTTFSS for 40m

I brought TTFSS set #7 to 40m and kept it in the electronic cabinet.

note that Q4 transistor has not been replaced back to PZT2907A yet. It's still GE82.

Q3 is now pzt3904, not PZT2222A.

 

  4511   Mon Apr 11 19:09:59 2011 SureshUpdateRF SystemInstalled low pass filters in the demod boards

 

As part of the RF system upgrade some of the demod boards in the lab were modfied.  The filter U5 (see the circuit schematic) was replaced. These changes are tabulated below.

 

Filters installed in the demod boards
Serial number Old name of the card New name of the card Filter installed Remarks
107 POY33 REFL33 SCLF-33+ R14=50Ohm
118 AP133, ASDD133 REFL55 SCLF-65  
114 PO199 REFL165 SCLF-190 R14=50Ohm
120 PO133 POP110 SCLF-135  
123 SP133 POP55 SCLF-65+ AT1 removed, R14=50Ohm
122 SP199, REFLDD199 AS165 SCLF-190  
121 SP166, REFL16 POP11 SCLF-10.7  
116 AP199 199 MHz POP165 SCLF-190  
126 AS166 33.3 MHz POX11 SCLF-10.7  
119 POX 33.3 MHz POY11 SCLF-10.7  
021 24.5 MHz (LLO) REFL11 SCLF-10.7  
020 24.5 MHz SCLF-45 POP22 SCLF-21.4  
022 24.5 MHz SCLF-45 AS11 with amp SCLF-10.7  
029 24.5          SCLF-f5 AS55 with amp SCLF-65  

 

Next, I and Q phase has to be checked for orthogonality. And noise levels of the cards have to measured.

 

 

 

  4510   Mon Apr 11 14:17:22 2011 josephb, jamieUpdateCDSFrame wiper script installed

[Joe, Jamie, Alex]

Fixes:

I asked Alex which cron to use (dcron? frcron?).  He promptly did the following:

emerge dcron

rc-update add dcron default

Copied the wiper.pl script from LLO to /opt/rtcds/caltech/c1/target/fb/

At that point, I modified wiper.pl script to reduce to 95% instead of 99.7%.

I added controls to the cron group on fb:

sudo gpasswd -a controls cron

I then added the wiper.pl to the crontab as the following line using crontab -e.

0 6 * * *       /opt/rtcds/caltech/c1/target/fb/wiper.pl --delete &> /opt/rtcds/caltech/c1/target/fb/wiper.log

Notes:

Note, placing backups on the /frames raid array will break this script, because it compares the amount in the /frames/full/, /frames/trends/minutes, and /frames/trends/seconds to the total capacity. 

Apparently, we had backups from September 27th, 2010 and March 22nd, 2011.  These would have broken the script in any case. 

We are currently removing these backups, as they are redundant data, and we have rsync'd backups of the frames and trends.  We should now have approximately twice the lookback of full frames.

  4509   Mon Apr 11 13:30:04 2011 josephb, JamieUpdateCDSNo Wiper script - Frames full over weekend

Problem:

The daqd process was dying every minute or so when it couldn't write frame.  This was slowing down the network by writing a 2.9G core dump over NFS every minute or so. (In /opt/rtcds/caltech/c1/target/fb/).

The problem was /frames/ was 100% full.

Apparently, when we switched the fb over to Gentoo, we forgot to install crontab and a wiper script.

Solution:

We will install crontab and get the wiper script installed.

  4508   Mon Apr 11 11:34:05 2011 KojiUpdateElectronicsNew Green PD test1

Quote:

 Ooh. Can you explain the purpose of the resistors which are connected to the (+) inputs? It looks like some real electronics ninjitsu.

51 Ohm for CLC409

The datasheet of CLC409 uses 25Ohm there. This is to cancel the input bias current of the two inputs of the opamp.

The source impedance (series) of SGD444 is 50Ohm. So I used 50Ohm for the + input shunting.

However, I could probably use anything between 0-50Ohm as the datasheet itself tells that the bias currents are
not related between the two inputs. In addition, I am not sure how much the real series resistance of the PD is.

1kOhm for OP27

This resister is to ensure the (+) input to have a high impedance at high frequencies.

As far as OP27 is behaving as an ideal opamp, the (+) input has a high impedance.
Also if the inductor behaves as the ideal inductor, no photocurrent comes to the AF path.

However, if both of the op27 and the inductor show similar impedances to the RF transimpedance of 240Ohm,
the AF path absorbs some photocurrent and affects the RF transimpedance of the RF output.

We know that the inductor has a self resonance where the shunt capacitance take over the impedance of the coil.
Above that frequency, the inductor is no longer the inductor. The self resonant freq of this inductor is ~300MHz. It is OK, but not
too far from the freq of interest if we like to see clear cut off at around f>100MHz.
Also OP27 is an AF amplifier and I had no confidence about the input impedance of the OP27 at 100~300MHz.

If I put 1k in the (+) input of the OP27, I can ensure the entire AF path has the impedance of ~1k (at least 500Ohm even when L and OP27 are shorted).
I think the chip resister easily works as a resister up to 1GHz.

Attachment 1: SGD444A_240_test.png
SGD444A_240_test.png
  4507   Mon Apr 11 09:49:53 2011 ranaUpdateElectronicsNew Green PD test1

 Ooh. Can you explain the purpose of the resistors which are connected to the (+) inputs? It looks like some real electronics ninjitsu.

  4506   Sun Apr 10 19:14:08 2011 KojiUpdateGreen LockingNew Green PD test1

I started to modify another green PD set.

It so far has the transimpedance of 240 Ohm on CLC409 for the RF output.

It shows the BB output upto ~100MHz.
The measurement shows the transimpedenca of ~90Ohm which is ~25% smaller than the expected gain of 120Ohm.
It is calibrated based on the transimpedances of Newfocus 1611 (10kOhm and 700Ohm for AF and RF).

The next step is to change the transimpedance resister to 2k and replace the PD to S3399 Si PD, which has the diameter of 3mm.
Then, the noise level will be measured. (and replace the RF opamp if necessary)

 

Attachment 1: SGD444A_240.png
SGD444A_240.png
Attachment 2: SGD444A_240_test.png
SGD444A_240_test.png
  4505   Fri Apr 8 20:43:46 2011 kiwamuSummaryIOORF combiner + resonant box : impedance and reflection coefficient

 The input impedance of the resonant box was measured when an RF combiner was attached to the box.

Indeed the combiner makes the impedance more 50 Ohm and reduces the reflection.

 

**** measurement conditions ****

* The output of box, where the EOM will be connected,  was open so that the box tries resonating with a parasitic capacitor instead of the real EOM.

* ZFSC-3-13, a 3-way combiner from mini circuit, was used.

* The S-port of the combiner was directly attached to the box with a short connector (~ 30 mm).

* Port 1 and 2 are terminated by 50 Ohm.

* The input impedance was measured on port 3 with AG4395A net work analyzer.

* Reflection coefficient 'Gamma' were calculated from the measured impedance 'Z' by using an equation Gamma = (50-Z)/(50+Z).

imp_refl.png

The resonances are found at 11, 29 and 73 MHz (55 MHz resonance was shifted to 73 MHz because of no EOM).

Note that the resonances are at frequencies where the notches appear in the reflection coefficient plot.

Don't be confused by a peak at 70 MHz in the impedance. This is an extra resonance due to a leakage inductance from the transformer in the circuit.

Quote: from #4504

An RF combiner should be included in the triple resonant box because it eases impedance mismatching and hence lowers undesired RF reflections.

  4504   Fri Apr 8 19:43:03 2011 kiwamuSummaryIOORF combiner eases impedance mismatching

An RF combiner should be included in the triple resonant box because it eases impedance mismatching and hence lowers undesired RF reflections.

Therefore we should use three cables to send the RF signals to the box and then combine them in the box.

 

(RF combiner)

 With proper terminations an RF combiner shows 50 Ohm input impedance.

But it still shows nearly 50 Ohm input impedance even if the source port is not properly terminated (i.e. non 50 Ohm termination).

This means any bad impedance mismatching on the source port can be somewhat brought close to 50 Ohm by a combiner.

  The amount of deviation from 50 Ohm in the input impedance depends on the circuit configuration of  the combiner as well as the termination impedance.

For example a resistive 3-way splitter shows 40 Ohm when the source port is shorten and the other ports are terminated with 50 Ohm.

Also it shows 62.5 Ohm when the source port is open and the other ports are terminated with 50 Ohm.

In this way an RF combiner eases  impedance mismatching on the source port.

 

(RF signal transfer at the 40m)

 According to the prototype test of the resonant box it will most likely have a non-50 Ohm input impedance at each modulation freqeucy.

If we install the resonant box apart from the combiner it will create RF reflections due to the mismatch (Case 1 in the diagram below)

The reflection creates standing waves which may excite higher harmonics and in the worst case it damages the RF sources.

 To reduce such a reflection one thing we can do is to install the combiner as a part of the resonant box (Case 2).

It will reduce the amount of the mismatching in the input impedance of the resonant circuit and results less reflections.

A rule we should remember is that a cable always needs to be impedance matched.

 

modulationBox.png

  4503   Fri Apr 8 01:05:45 2011 SureshUpdateRF SystemRF Source Harmonics

 

 The measured power levels of the RF source harmonics are given below:

 

 

We are considering inclusion of bandpass filters centered on 11 and 55 MHz  to suppress the harmonics and meet the requirements specified in Alberto's thesis (page 88).

 

Attachment 1: RF_Source_Harmonics_Sheet1.pdf
RF_Source_Harmonics_Sheet1.pdf RF_Source_Harmonics_Sheet1.pdf RF_Source_Harmonics_Sheet1.pdf RF_Source_Harmonics_Sheet1.pdf
  4502   Thu Apr 7 21:58:57 2011 AidanSummaryGreen LockingBeat note amplitude

Having convinced myself that the green Hartmut PD is giving an acceptable response at RF frequencies I decided to double-check the beatnote at IR (fiber transmission from the X-end beating with the PSL). This took a while because I had to realign the beam into the fiber at the X-end (I had a PD monitoring the output from the fiber on the PSL table and 40m of BNC cable giving me the signal from it at the X-end).

Eventually, I managed to get a beatnote on the PD. At first there was no signal at the temperature calculated using Koji and Suresh's calibration, but it turned out that the mode-overlap wasn't good enough on the PD. Now I can clearly see beats between a couple of modes, one of which is much stronger than the other. I think we should use a frequency discriminator on the output from the IR PD to servo the end laser and keep the strong beat note within <100MHz of DC.

 

  4501   Thu Apr 7 19:28:02 2011 KojiSummaryGreen Locking(In)sanity check of Green PD - some inconsistencies

Responsivity of SGD-444A

Quote:

For instance, the PDA100A Si detector from Thorlabs has a responsivity of 0.35-0.4A/W at 1064nm.

 

Attachment 1: SGD-444A.png
SGD-444A.png
  4500   Thu Apr 7 16:09:17 2011 AidanSummaryGreen Locking(In)sanity check of Green PD - some inconsistencies

I think I had underestimated the responsivity of the Silicon PD at 1064nm. The previous value was based on a rough search online for the responsivity of Silicon (I couldn't find the product number of the actual PD we are using). For instance, the PDA100A Si detector from Thorlabs has a responsivity of 0.35-0.4A/W at 1064nm. 

If we calculate the responsivity of the Hartmut PD from the measurements I made today (input power = 0.300mW, output voltage = 5.56mV, effective transimpedance = 80 Ohms), then the responsivity at 1064nm is 0.23 A/W which is not an unreasonable number given the response of the Thorlabs detector.

Quote:

Measurement Value
Responsivity of Silicon PD at 1064nm 0.02 A/W (estimate)
Responsivity of 1611 New Focus PD at 1064nm ~0.8 A/W
   

There is one other troubling point: using the estimate of responsivity on the Harmut PD * incident power * transimpedance at DC = (0.02A/W) * (0.28mW) * (80 V/A) = 0.45 mV.

But the measured DC voltage is 6.5mV = inconsistent.

 

  4499   Thu Apr 7 13:14:23 2011 josephbUpdateCDSProposed plan for ITMX/ITMY control switch

Problem:

The controls (fast and slow both) think ITMX is ITMY and ITMY is ITMX.

Solution:

After some poking around today, I have convinced myself it is sufficient to simply swap all instances of ITMX for ITMY in the C1_SUS-AUX1_ITMX.db  file, and then rename it to C1_SUS-AUX1_ITMY.db (after having moved the original C1_SUS-AUX1_ITMY.db to a temporary holding file).

A similar process is then applied to the original C1_SUS-AUX1_ITMY.db file.  These files live in /cvs/cds/caltech/target/c1susaux.  This will fix all the slow controls.

To fix the fast controls, we'll modify the c1sus.mdl file located in /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/simLink/ so that the ITMX suspension name is changed to ITMY and vice versa.  We'll also need to clean up some of the labeling

At Kiwamu and Bryan's request, this will either be done tomorrow morning or on Monday.

So the steps in order are:

1) cd /cvs/cds/caltech/target/c1susaux

2) mv C1_SUS-AUX1_ITMX.db C1_SUS-AUX1_ITMX.db.20110408

3) mv C1_SUS-AUX1_ITMY.db C1_SUS-AUX1_ITMY.db.20110408

4) sed 's/ITMX/ITMY/g' C1_SUS-AUX1_ITMX.db.20110408 > C1_SUS-AUX1_ITMY.db

5) sed 's/ITMY/ITMX/g' C1_SUS-AUX1_ITMY.db.20110408 > C1_SUS-AUX1_ITMX.db

6) models

7) matlab

8) Modify c1sus model to swap ITMX and ITMY names while preserving wiring from ADCs/DACs/BO to and from those blocks.

9) code; make c1sus; make install-c1sus

10) Disable all watchdogs

11) Restart the c1susaux computer and the c1sus computer

 

  4498   Thu Apr 7 13:12:23 2011 KojiHowToVIDEOVideo switching tip

Long time ago, I looked at the manual of the video switcher.
http://media.extron.com/download/files/userman/Plus_Ultra_MAV_C.pdf
Here is the summary. This will be the basic of the more sophisticated switching program which may have GUI.

In principle, you can manually control the matrix via telnet. At the console machines, you can connect to the matrix using telnet

telnet 192.168.113.92

This opens TCP/IP port 23 of the specified machine. You will receive some messages.
Then type some command like:
--------------------

  • 1*2!       (connect input#1 to output#2)
  • 1,           (save the current setting into preset1)
  • 1.           (restore the setting from preset1)

--------------------

Basicaly that's all. There are many other features but I don't think we need them.

We can create a simple program with any of the language as any of the language has the capability of the TCP/IP connection.
e.g. C, Perl, Python. Tcl/Tk
Any of them are fine.

Now what we have to think about is how to implement the interface in the epics screen (or whatever).
It needs some investigation how the people is thinking as the ideal interface.
But, first of all, you should make the above three operations available as a simple UNIX command like:

videoswitch -i 192.168.113.92 1 2
videoswitch -i 192.168.113.92 -store 1
videoswitch -i 192.168.113.92 -recall 1
(There is no such command yet. These are showing what it should be!)

This can be done by a single day work and our life will be much better.

  4497   Thu Apr 7 11:51:13 2011 steveSummarySAFETYnew crane operator inaugurated

Quote:

Mike Caton of Konecranes inspected and loadtested all 3  of the 40m cranes at max reach trolley positions with 1 ton.

 Konecrane representative gave crane operator training in the 40m. Koji has become a qualified, trained crane operator of the 40m lab.

Attachment 1: P1070535.JPG
P1070535.JPG
  4496   Thu Apr 7 11:38:56 2011 steveUpdatePSLenclosure windows on the east side of the PSL

The PSL enclosure now have 4 windows on each side. The bottom rail guides on the east side will be replaced by  one  U-channel for smoother, more gentle sliding.

Door position indicator- interlock switches are not wired yet.

Attachment 1: P1070538.JPG
P1070538.JPG
  4495   Wed Apr 6 22:13:24 2011 BryanConfigurationGreen LockingResonating green light!

Every so often things just work out. You do the calculations, you put the lenses on the bench, you manually adjust the pointing and fiddle with the lenses a bit, you get massive chunks of assistance from Kiwamu to get the alignment controls and monitors set up and after quite a bit of fiddling and tweaking the cavity mirror alignment you might get some nice TEM_00 -like shapes showing up on your Y-arm video monitors.

So. We have resonating green light in the Y-arm. The beam is horribly off-axis and the mode-matching, while close enough to give decent looking spots, has in no way been optimised yet. Things to do tomorrow - fix the off-cavity-axis problem and tweak up the mode-matching... then start looking at the locking...

  4494   Wed Apr 6 19:36:32 2011 AidanSummaryGreen Locking(In)sanity check of Green PD - some inconsistencies

I moved the Hartmut Green PD to the Jenne laser bench to try to determine if the response at RF was reasonable or somehow very much smaller than it should be. It was set up as shown in the attached diagram. The first pass at this was by comparing the ratio of the RF photocurrent of the green PD to the RF photocurrent of the New Focus 1611 InGaAs PD. That ratio (at a sufficiently low frequency) should be the same as the ratio the DC photocurrents of the two PDs.

Using the network analyzer I measured the ratio of the voltages of the two RF signals (and then scaled each of these by the respective transimpedances of the PDs: 700 Ohms for the 1611 and 240 Ohms for the Harmut PD). The resulting ratio is shown in the attached plot.

I measured the DC voltages from each PD and scaled those by the transimpedances to get the photocurrent (10 kOhm for the 1611 and 80 Ohm effective for the Harmut PD). The ratio of the DC photocurrents was 0.37. This is roughly 3x the ratio of the RF photocurrents at 500kHz (=0.14). This discrepancy is uncomfortably large.

 The full set of measurements is given in the table below:

Measurement Value
DC voltage from Hartmut PD 6.5mV (checked by turning laser on and off and measuring the difference)
DC voltage from 1611 InGaAs PD 2.20V
Transimpedance of Harmut PD at DC 80 Ohm (effective)
Transimpedance of Harmut PD at RF 240 Ohm
Transimpedance of 1611 InGaAs at DC 10 KOhm
Transimpedance of 1611 InGaAs at RF 700 Ohm
Incident Power on Hartmut PD (100% on PD area) 0.28mW (measured by Ophir power meter)
Incident Power on 1611 InGaAs (<100% on PD area) 0.64mW
Responsivity of Silicon PD at 1064nm 0.02 A/W (estimate)
Responsivity of 1611 New Focus PD at 1064nm ~0.8 A/W
   

There is one other troubling point: using the estimate of responsivity on the Harmut PD * incident power * transimpedance at DC = (0.02A/W) * (0.28mW) * (80 V/A) = 0.45 mV.

But the measured DC voltage is 6.5mV = inconsistent.

Attachment 1: PD_measurement.png
PD_measurement.png
Attachment 2: plot_PD_RF_ratios.pdf
plot_PD_RF_ratios.pdf
  4493   Wed Apr 6 18:55:49 2011 Jamie, LarisaConfigurationLSCmajor AP table cleanup

We ripped out all of the old AS, PLL, and REFL paths, green, orange, and cyan respectively on the old AP table layout photo:

  • AS (green): had already been re-purposed by putting a ThorLabs diode right after the first steering mirror.   Everything downstream of that has been removed.
  • PLL (orange): everything removed.
  • REFL (cyan): CCD was left in place, so everything upstream of that was not touched.  Everything else was removed, including all of the REFL detectors.
  • OMCT (purple): previously removed
  • OMCR (blue): left in place, but the diode and CCD are not connected (found that way).
  • MCT (magenta): previously removed.
  • IMRC (red): untouched

All optics and components were moved to the very south end of the SP table.

We also removed all spurious cables from the table top, and from underneath, as well as pulled out no-longer-needed power supplies.

  4492   Wed Apr 6 16:02:07 2011 Larisa ThorneUpdateElectronicsCable laying...continued

[Steve, Kiwamu, Larisa]

 

Having finished laying new cable last week, we moved on to connecting those on PSL table and AP table.

Cables connected:

--RCR, RCT, PMCR (all three are blue)

--OMCR (blue cable, ***now has a camera***), PMCT, IMCR, REFL, AS (white cable), OMCT (***now has camera***)

 

Unless otherwise noted, the cables are black on the AP table. Also on the AP table: cables were connected directly to the power source.

The wiki has been updated accordingly.

 

Steve noted that MC2T and POP cameras are not there.

 

 

  4491   Wed Apr 6 02:41:01 2011 kiwamuUpdateGreen Lockingnoise budget : some more noise

It turned out that the dark noise from the beat PD and the shot noise on the beat PD was overestimated.

So I corrected them in the plot of the last noise budget (#4482).

Additionally I added the end laser error signal in the plot. Here is the latest plot.

noise_budget.png

 The end laser error spectrum is big enough to cover most of the frequency range.
 (although it was taken at a different time from the other curves.)

Quote from #4482

According to the plot what we should do are :

  * More accurate PD noise measurement

  * More accurate shot noise estimation

  4490   Tue Apr 5 21:20:11 2011 KojiSummaryGreen LockingX-arm cavity locked with LB1005 servo box

Last Thursday, I tested Newport Servo Controller LB1005 with the X_arm green PDH servo.

The setup and the settings I could lock the arm is depicted in the attached figure.
To lock the cavity, follow the steps below

1) Toggle the switch to the "lower" position. This disengages the servo and reset the integrator.

2) Toggle the switch to the "middle" position. The zero freq is set to the "PI corner" freq. At the low freq the gain is limited
at the value of "LF Gain Limit". This gives us a single pole at the low freq.

3) Once the lock is acquired, toggle the switch to the "upper" position. This moves the pole freq to DC, resulting in the complete integration of the signal at the low frequency.

I measured the openloop transfer function (attachment 2). The amp is quite fast and exhibits almost no phase delay upto 100kHz.
The UGF was 10kHz with the phase mergin of ~45deg. I had to tune the input offset carefully to stay at the center of the resonance.

Attachment 1: servo.png
servo.png
Attachment 2: OLTF.pdf
OLTF.pdf
  4489   Tue Apr 5 19:54:39 2011 KojiSummaryGreen LockingHamamatsu S3399 test

Since last Friday I have been testing the broadband RF photodetector in order to figure out the capability of S3399 with the similar circuit as Matt's BBPD
We also like to figure out if it has sufficient performance for the 40m green locking.

The circuit diagram is shown in the first attachment. The RF amplifier is attached at the diode while the reverse bias voltage is applied at the other side of the diode. The amplifier's input impedance is used as the transimpedance resister. Note that the bandwidth of this configuration is limited by the RC filter that consists of the junction capacitance of the diode, the series resistance of the diode, and the transimpedance resister. This cut off freq is in general lower than that cut off obtained with the usual transimpedance amplifier which has the readout resister at the feedback path of the opamp.

The transfer function of the PD is measured using Jenne's laser. At the reverse bias voltage of 30V, the -3dB bandwidth of 178MHz was obtained. This is quite high bandwidth for the most of the applications at the 40m.

Because of the low transimpedance the low-noise level of the RF amplifier is very crucial. Recently we can obtain an ultra low noise RF amplifier like Teledyne Cougar AC688 which has the NF of 0.9dB with the bandwidth between 10MHz - 600MHz. Next step will be to obtain this kind of amplifier to test the noise performance.

 

 

Attachment 1: S3399_test_110405.png
S3399_test_110405.png
Attachment 2: S3399_test_110405.pdf
S3399_test_110405.pdf
  4488   Tue Apr 5 17:31:59 2011 steveSummaryGeneralnew laser pointers

Quote:

Just for a record. We got 4 new laser pointers (2 greens, 1 blue, and 1 green and red combination). Don't lose them.

They reside in a bucket on the SP table, where IR viewers and sensor cards also reside.

DSC_2883_ss.jpg

 Low power green-red laser pointers are in. High power green, red and blue pointers are confiscated.

Attachment 1: P1070530.JPG
P1070530.JPG
  4487   Tue Apr 5 17:04:36 2011 steveSummarySAFETYcranes inspected and load tested

Mike Caton of Konecranes inspected and loadtested all 3  of the 40m cranes at max reach trolley positions with 1 ton.

Attachment 1: P1070522.JPG
P1070522.JPG
Attachment 2: P1070532.JPG
P1070532.JPG
  4486   Mon Apr 4 18:58:44 2011 BryanConfigurationGreen LockingA beam of purest green

We now have green light at the Y end. 

The set-up (with careful instructions from Kiwamu) - setting up with 100mW of IR into the oven.

Input IR power = 100mW measured.

 

Output green power = 0.11mW

(after using 2 IR mirrors to dump IR light before the power meter so losing a bit of green there light too)

 

And it's pretty circular-looking too. Think there might be a bit more efficiency to be gained near the edges of the crystal with internal reflections and suchlike things but that gives us an UGLY looking beam.  Note - the polarisation is wrong for the crystal orientation so used a lambda/2 plate to get best green  power out.

 

Efficiency is therefore 0.11/100 = 0.0011 (0.11%) at 100mW input power.

 

Temperature of the oven seems to be around 35.5degC for optimal conversion.

Took a picture. Ta-dah! Green light, and lots more where that came from! Well... about 3x more IR available anyway.

 

P4040042.JPG

 

 

  4485   Mon Apr 4 14:20:32 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Last bit of oven matching for now.

 

I moved the lens before the oven position back along the beam path by about 1cm - waist should be just above position 9 in this case. Note - due to power-findings from previous time I'm maximising the power into the head to reduce the effect of offsets.

 

From position 9:

Position A1_13.5%_width A2_13.5%_width

(mm) (um mean) (um mean)

-1 121.1 123.6

0 112.5 113.8

1 106.4 106.1

2 102.9 103.4

3 103.6 103.6

4 106.6 107.4

5 111.8 112.5

6 118.2 120.1

7 126.3 128.8

8 134.4 137.1

9 143.8 146.5

10 152.8 156.1

11 163.8 167.1

12 175.1 176.4

13 186.5 187.0

14 197.1 198.4

15 210.3 208.9

16 223.5 218.7

17 237.3 231.0

18 250.2 243.9

19 262.8 255.4

20 274.7 269.0

21 290.4 282.3

22 304.3 295.5

23 316.7 303.1

 

Note - had to reduce power due to peak saturation at 15mm - don't think scale changed, but be aware just in case. And saturated again at 11. And again at 7. A little bit of power adjustment each time to make sure the Beamscan head wasn't saturating. Running the fit gives...

 

Waist_Fits_from_laser.pngWaist_Fits_Bench_Position.png

 

OK. The fit is reasonably good. Residuals around the area of interest (with one exception) are <+/- 2um and the waists are 47.5um (vertical) and 50.0um (horizontal) at a position of 9.09 on the bench. And the details of the fitting output are given below.

 

-=-=-=-=-=-=-=-=-=-=-=-

Vertical Fit

 

cf_ =

 

     General model:

       cf_(x) = w_o.*sqrt(1 + (((x-z_o)*1064e-9)./(pi*w_o.^2)).^2)+c

     Coefficients (with 95% confidence bounds):

       c =   5.137e-06  (4.578e-06, 5.696e-06)

       w_o =   4.752e-05  (4.711e-05, 4.793e-05)

       z_o =        1.04  (1.039, 1.04)

 

 

cfgood_ = 

 

           sse: 1.0699e-11

       rsquare: 0.9996

           dfe: 22

    adjrsquare: 0.9996

          rmse: 6.9738e-07

 

-=-=-=-=-=-=-=-=-=-=-=-

Horizontal Fit

 

cf_ =

 

     General model:

       cf_(x) = w_o.*sqrt(1 + (((x-z_o)*1064e-9)./(pi*w_o.^2)).^2)+c

     Coefficients (with 95% confidence bounds):

       c =    3.81e-06  (2.452e-06, 5.168e-06)

       w_o =   5.006e-05  (4.909e-05, 5.102e-05)

       z_o =        1.04  (1.04, 1.04)

 

 

cfgood_ = 

 

           sse: 4.6073e-11

       rsquare: 0.9983

           dfe: 22

    adjrsquare: 0.9981

          rmse: 1.4471e-06

 

 

 

  4484   Mon Apr 4 11:52:13 2011 JenneUpdatePEMSTS2s unpacked

I unpacked the STS2 seismometers that we borrowed from LLO.  They are sitting underneath the Xarm, in the middle of the mode cleaner, near the other seismometer stuff. 

  4483   Fri Apr 1 23:49:24 2011 kiwamuSummaryGreen Lockingtwo states in green beat-note

According to the measurement done by Aidan and me, there are two beat-note state.

One gave us a small beat signal and the other gave us a bigger signal by approximately 20 dB.

 

 A possible reason for this phenomenon is that the end laser is operating at a special temperature that somehow drives the laser with two different modes at the same time.

So that it permits the laser sometimes locked with one of the two modes and sometimes with the other mode.

For the first step we will change the temperature such that the laser can run with a single stable mode.

Then for investigating it we will put a scanning cavity on the X end table to see if it really exhibits a two modes or not.

Quote from #4472

The attached table shows the amplitude of the green beat note when the end laser was in various states. We can increase the beat note amplitude dramatically by switching to a different states.

  4482   Fri Apr 1 23:05:58 2011 kiwamuUpdateGreen Lockingnoise budget

I made a coarse noise budget in order to decide our next actions for the X arm green locking.

So be careful, this is not an accurate noise budget !

 Some data are just coming from rough estimations and some data are not well calibrated.

noise_budget.png

 Assuming all the noise are not so terribly off from the true values, the noise at high frequency is limited by the dark noise of the PD or it already reaches to the IR inloop signal.

The noise at low frequency is dominated by the intensity noise from the transmitted green light although we thought it has been eliminated by the comparator.

In any case I will gradually make this noise budget more accurate  by collecting some data and calibrating them.

 

According to the plot what we should do are :

  * More accurate PD noise measurement

  * More accurate shot noise estimation

  * Searching for a cause of the small beat signal (see here) because a bigger beat signal lowers the PD noise.

  * Investigation of the Intensity noise

  4481   Fri Apr 1 18:54:41 2011 BryanConfigurationGreen LockingY end doubling oven

The doubling oven is now ready to go for the Y arm. The PPKTP crystal is mounted in the oven:

P4010036.JPG

Note - the crystal isn't as badly misaligned as it looks in this photo. It's just an odd perspective shot. I then closed it up and checked to make sure the IR beam on the Y bench passes through the crystal. It does. Just need to tweak the waist size/position a bit and then we can actually double some frequencies!

P4010041.JPG

  4480   Thu Mar 31 20:46:11 2011 AidanSummaryGreen LockingGreen beat note PD DC response

I measured the DC response of the Green PD


Power into PD at DC (green laser pointer) = 285 uW
Voltage out of PD = 552mV/(100x SR560gain) = 5.52mV
Photocurrent = 5.52mV/(241 Ohms)*3 = 68.7uA
Responsivity = 68.7/285 = 0.24 A/W

Therefore, since the responsivity is in the correct range for a Silicon PD at 532nm, the DC output is giving us sensible response to an input signal.


But, there is a 2.12MHz, 328mV oscillation on the DC output irrespective of the incident power.
 

  4479   Thu Mar 31 20:37:10 2011 AidanSummaryGreen LockingRF amplitude source

 I gutted one of the $2 red laser pointers to build a laser source whose amplitude we could modulate at RF frequencies. Basically, I cut off the bulk of the housing from the pointer and soldered a BNC connection into the two terminals that the 2x 1.5V batteries were connected to. When I applied 3V across this BNC connector the diode still worked. So far so good.

Next I added a bias tee to the input. I put 3V across the DC input of the bias tee and added a -3dBm signal into the RF port of the tee. The laser beam was incident on a PDA100A (bandwidth of 1.7MHz) and, sure enough, Kiwamu and I could see a flat response in the amplitude at a given drive frequency out to around 1.7MHz.

We should check the response on a faster PD to see how fast the laser diode is, but we should be able to use this now to check the RF response of the green beat note PD. 

TO DO:

1. Add some capacitors across the DC input of the bias tee.

2. Do something about the switch on the laser diode.

3. Attach some labels to the laser that specify what is the required DC voltage and the maximum acceptable RF modulation amplitude.

Attachment 1: P1000543.jpg
P1000543.jpg
Attachment 2: P1000544.jpg
P1000544.jpg
Attachment 3: P1000545.jpg
P1000545.jpg
  4478   Thu Mar 31 19:58:11 2011 kiwamuUpdateCDSc1iscex crashed

After I did several things to add new DAQ channels on c1iscex it suddenly became out of network. Maybe crashed.

Then c1iscex didn't respond to a ping and all the epics values associated with c1iscex became not accessible.

I physically shut it down by pushing the reset button. Then it came back and is now running fine.

 


(how I broke it)

Since activateDAQ.py has screwed up the 'ini' files including C1SCX.ini, I was not able to add a channel to C1SCX.ini by the usual daqconfig GUI.

So I started editing it in a manual way with an editor and changed some sentences to that shown below

  [C1:GCX-ERR_MON_IN1_DAQ]
  acquire=1
  chnnum=10004
  datarate=2048
  datatype=4
  [C1:GCX-GRN_REFL_DC_IN1_DAQ]
  acquire=1
  chnnum=10007
  datarate=2048
  datatype=4
  [C1:GCX-SLOW_SERVO1_IN1_DAQ]
  acquire=1
  chnnum=10010
  datarate=2048
  datatype=4

Then I rebooted fb to reflect the new DAQ channels.

After that I looked at the C1_FE_STATUS.adl screen and found some indicator lights were red.

So I pushed "Diag reset" button and "DAQ Reload" button on the C1SCX_GDS_TP.adl screen and then c1iscex died.

After the reboot the new DAQ channels looked acquired happily.

This is my second time to crash a front end machine (see this entry)

  4477   Thu Mar 31 15:23:14 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Quote:

3. Zoomed-in plot for the SHG crystal show no astigmatism. However, the zoomed out plot shows some astigmatism.

How consistent are they? ==> Interested in seeing the fit including the zoomed out measurements.

Right. Fitting to the data. Zoomed out plots first. I used the general equation f(x) = w_o.*sqrt(1 + (((x-z_o)*1064e-9)./(pi*w_o.^2)).^2)+c for each fit which is basically just the Gaussian beam width parameter calculation but with an extra offset parameter 'c'

Vertical fit for zoomed out data:

Coefficients (with 95% confidence bounds):

       c =   7.542e-06  (5.161e-06, 9.923e-06)

       w_o =   3.831e-05  (3.797e-05, 3.866e-05)

       z_o =       1.045  (1.045, 1.046)

 

Goodness of fit:

  SSE: 1.236e-09

  R-square: 0.9994

 
Horizontal fit for zoomed out data:
 

Coefficients (with 95% confidence bounds):

       c =   1.083e-05  (9.701e-06, 1.195e-05)

       w_o =   4.523e-05  (4.5e-05, 4.546e-05)

       z_o =       1.046  (1.046, 1.046)

 

Goodness of fit:

  SSE: 2.884e-10

  R-square: 0.9998

  Adjusted R-square: 0.9998

  RMSE: 2.956e-06

 

Zoomed_out_fitting01.png

-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-

 

OK. Looking at the plots and residuals for this, the deviation of the fit around the waist position, and in fact all over, looks to be of the order 10um. A bit large but is it real? Both w_o values are a bit lower than the 50um we'd like, but… let's check using only the zoomed in data -  hopefully more consistent since it was all taken with the same power setting.

 

 

Vertical data fit using only the zoomed in data:

 

Coefficients (with 95% confidence bounds):

       c =   1.023e-05  (9.487e-06, 1.098e-05)

       w_o =   4.313e-05  (4.252e-05, 4.374e-05)

       z_o =       1.046  (1.046, 1.046)

 

Goodness of fit:

  SSE: 9.583e-11

  R-square: 0.997

 

Horizontal data fit using only the zoomed in data:

 

Coefficients (with 95% confidence bounds):

       c =   1.031e-05  (9.418e-06, 1.121e-05)

       w_o =    4.41e-05  (4.332e-05, 4.489e-05)

       z_o =       1.046  (1.046, 1.046)

 

Goodness of fit:

  SSE: 1.434e-10

  R-square: 0.9951

 

-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-

Zoomed_in_fitting01.png

 

The waists are both fairly similar this time 43.13um and 44.1um and the offsets are similar too  - residuals are only spread by about 4um this time.

 

I'm inclined to trust the zoomed in measurement more due to the fact that all the data was obtained under the same conditions, but either way, the fitted waist is a bit smaller than the 50um we'd like to see. Think it's worthwhile moving the 62.9mm lens back along the bench by about 3/4 -> 1cm to increase the waist size.

 

 

 

 

 

  4476   Thu Mar 31 14:10:00 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Quote:

 I went through the entries.

1. Give us a photo of the day. i.e. Faraday, tilted lens, etc...

2. After all, where did you put the faraday in the plot of the entry 4466?

3. Zoomed-in plot for the SHG crystal show no astigmatism. However, the zoomed out plot shows some astigmatism.
How consistent are they? ==> Interested in seeing the fit including the zoomed out measurements.

 OK. Taking these completely out of order in the easiest first...

2. The FI is between positions 27.75 and 32 on the bench - i.e. this is where the input and output apertures are. (corresponds to between 0.58 and 0.46 on the scale of those two plotsand just before both the vertical and horizontal waists) At these points the beam radius is around 400um and below, and the aperture of the Faraday is 4.8mm (diameter).

1. Photos...

Laser set up - note the odd angles of the mirrors. This is where we're losing a goodly chunk of the light. If need be we could set it up with an extra mirror and send the light round a square to provide alignment control AND reduce optical power loss...

P3310028.JPG

 

Faraday and angled lens - note that the lens angle is close to 45 degrees. In principle this could be replaced with an appropriate cylindrical lens, but as long as there's enough light passing through to the oven I think we're OK.

P3310029.JPG

3. Fitting... coming soon once I work out what it's actually telling me. Though I hasten to point out that the latter points were taken with a different laser power setting and might well be larger than the actual beam width which would lead to astigmatic behaviour.

  4475   Thu Mar 31 11:30:51 2011 steveConfigurationGeneralstrain relieved rf cables

I strain relieved RF cables labeled 33 MHZ LO and 166 MHZ to EOM at 1X2  This is a temporary setup for the 11 MHZ

The coax N  bulkheads connectors are mounted on the plastic front panel now.

Attachment 1: P1070491.JPG
P1070491.JPG
  4474   Thu Mar 31 08:31:44 2011 SureshUpdateVIDEOCable laying...continued

The video work has crossed a milestone.    

Kiwamu and Steve have shifted the three quads from the control room to the Video MUX rack (1Y1) and have wired them to the MUX.

The monitors in the control room have been repositioned and renumbered.  They are now connected directly to the MUX. 

Please see the new cable list for the input and output channels on the MUX.

As of today, all cables according the new plan are in place.  Their status   indicated on the wiki page above is not verified .  Please ignore that column for now, we will be updating that soon.

I shifted the MC1F/MC3F camera and the MC2F cameras onto the new cables.  Also connected the monitors at the BS chamber and end of the X arm to their respective cables.  I have removed the RG58B BNC (black) cables running from MC2 to BS and from ETMXF to the top of the Flow Bench. 

Some of the old video cables are still in place but are not used.   We might consider removing them to clear up the clutter. 

Some of the video cables in use are orange and if the lab's  cable color code is to be enforced these will have to be replaced with blue ones..

Some of the cables in use running from the MUX to the monitor in the control room are the white 50 Ohm variety.  There are also black RG59 Cables running the same way ( we have surplus cables in that path)  and we have to use those instead of the white ones. 

There are a number of tasks remaining:

a)  The inputs from the various existing cameras have to be verified. 

b) There are quite a few cameras which are yet to be installed.

c) The Outputs may not not be connected to their monitors.  That the monitors may still be connected to an old cable which is not connected to the MUX.  The new cable should be lying around close by.  So if you see a blank monitor please connect it to its new cable. 

d) The status column on the wiki page has to be updated.

e) Some of the currently in place may need to be replaced and some need to be removed.  We need to discuss our priorities and come up with a plan for that.

After checking everything we can certify that the video cabling system is complete.

I would like Joon Ho to take care of this verification+documenting process and declaring that the job is complete. 

 

Steve attached these two pictures.

Attachment 1: P1070489.JPG
P1070489.JPG
Attachment 2: P1070494.JPG
P1070494.JPG
ELOG V3.1.3-